# Deep Attention Driven Reinforcement Learning (DAD-RL) for Autonomous Vehicle Decision-Making in Dynamic Environment

### v1

@@ III-A Observation Space (Input) Preprocessing
The ego-vehicle can obtain historical information of its surrounding vehicles $H_{t}$, segmented Bird-Eye-View (BEV) context $C_{t}$, and an elaborate ego’s odometric state history $E_{t}$ as shown in Fig. 2. The whole observation space tensor is represented as $\mathcal{O}_{t}=\left[H_{t};E_{t};\mathcal{C}_{t}\right]$. The subscript $t$ attributes to the tensor’s value at time $t$. The BEV context is defined as $\mathcal{C}_{t_{o}}=\{D_{t};W_{t}\}$, where $D_{t}$ is drivable area map and $W_{t}$ waypoint map of size $128\times 128$. The historical surrounding vehicle data is defined as ${H}_{t}=\left[H^{1}_{t},...,H^{n}_{t}\right]$ for $n$ surrounding vehicles, with $H^{i_{sv}}_{t}=\left[\left(X^{i_{sv}}_{t-5k.\delta t},\phi^{i_{sv}}_{t-5k.\delta t},v^{i_{sv}}_{t-5k.\delta t},l^{i_{sv}}_{t-5k.\delta t}\right)\right]^{4}_{k=0}$ of vehicle $i$’s current and past timesteps spaced four simulation steps $\delta t$ apart. Here, for vehicle $i$, $X^{i_{sv}}_{t}=(x^{i_{sv}}_{t},y^{i_{sv}}_{t})$, $\phi^{i_{sv}}_{t}$, $v^{i_{sv}}_{t}$, and $l^{i_{sv}}_{t}$ are historical data of the relative position to the ego, the heading, speed of the vehicle and the lane it is in respectively. Observation $E_{t}=\left[\left(E^{(1)}_{t};E^{(2)}_{t}\right)\right]$ consists of various historical odometric values split into two vectors based on utility in the DAD-RL framework. The vector $E^{(1)}_{t}$ contains the same quantities as in $H^{i_{sv}}_{t}$ and $E^{(2)}_{t}=\left[\left(\omega_{t-5k.\delta t},\psi_{t-4k.\delta t},v^{ego}_{t-5k.\delta t},a^{ego}_{t-5k.\delta t},j_{t-5k.\delta t}\right)\right]^{4}_{k=0}$, a tuple having steering, yaw rate, linear velocity, linear acceleration, and linear jerk of the ego at time $t$ respectively. The following section explains how the spatiotemporal deep attention encoder processes this observation space. The encoder has two parts: a) Spatio-Temporal Attention Encoder (STAE) and b) Context Encoder (CE).
@@ IV-C Comparison Baselines
To thoroughly compare the proposed framework’s performance, the DAD-RL framework was evaluated against several baseline methods, and SOTA [11]. PPO: Proximal Policy Optimization stands out as a SOTA policy gradient approach with promising performance on robotic decision-making tasks. SAC: Soft Actor-Critic is another SOTA RL algorithm, an off-policy approach capitalizing on entropy maximization to assist agent training. DrQ: Data-regularized Q-learning employs a CNN-based augmentation to SAC, facilitating direct learning of a policy function from pixels. RDM: Rule-based Driver Model mimicking the driving behavior of neighboring vehicles within SMARTS simulations, offering a benchmark for evaluating the rule-based performance. DT: Decision Transformer is an innovative approach that reformulates the RL problem as a conditional sequence modeling task and employs a transformer to generate future actions from past states, actions, and rewards. SRT: Scene-Rep Transformer (SRT) is a transformer-based framework recently proposed in [11]. SRT employs two transformer architectures - a Multi-Stage Transformer (MST) for encoding the multi-modal scene input and another Sequential Latent Transformer (SLT) to instill predictive information into the latent state vector.

### v2

# Deep Attention Driven Reinforcement Learning (DAD-RL) for Autonomous Decision-Making in Dynamic Environment

@@ III-A Observation Space (Input) Preprocessing
The ego-vehicle can obtain historical information of its surrounding vehicles $H_{t}$, segmented Bird-Eye-View (BEV) context $C_{t}$, and an elaborate ego’s odometric state history $E_{t}$ as shown in Fig. 2. The whole observation space tensor is represented as $\mathcal{O}_{t}=\left[H_{t};E_{t};\mathcal{C}_{t}\right]$. The subscript $t$ attributes to the tensor’s value at time $t$. The BEV context is defined as $\mathcal{C}_{t_{o}}=\{D_{t};W_{t}\}$, where $D_{t}$ is drivable area map and $W_{t}$ waypoint map of size $128\times 128$. The historical surrounding vehicle data is defined as ${H}_{t}=\left[H^{1}_{t},...,H^{n}_{t}\right]$ for $n$ surrounding vehicles, with $H^{i_{sv}}_{t}=\left[\left(X^{i_{sv}}_{t-5k.\delta t},\phi^{i_{sv}}_{t-5k.\delta t},v^{i_{sv}}_{t-5k.\delta t},l^{i_{sv}}_{t-5k.\delta t}\right)\right]^{4}_{k=0}$ of vehicle $i$’s current and past timesteps spaced four simulation steps $\delta t$ apart. Here, for vehicle $i$, $X^{i_{sv}}_{t}=(x^{i_{sv}}_{t},y^{i_{sv}}_{t})$, $\phi^{i_{sv}}_{t}$, $v^{i_{sv}}_{t}$, and $l^{i_{sv}}_{t}$ are historical data of the relative position to the ego, the heading, speed of the vehicle and the lane it is in respectively. Observation $E_{t}=\left[\left(E^{(1)}_{t};E^{(2)}_{t}\right)\right]$ consists of various historical odometric values split into two vectors based on utility in the DAD-RL framework. The vector $E^{(1)}_{t}$ contains the same quantities as in $H^{i_{sv}}_{t}$ and $E^{(2)}_{t}=\left[\left(\omega_{t-5k.\delta t},\psi_{t-5k.\delta t},v^{ego}_{t-5k.\delta t},a^{ego}_{t-5k.\delta t},j_{t-5k.\delta t}\right)\right]^{4}_{k=0}$, a tuple having steering, yaw rate, linear velocity, linear acceleration, and linear jerk of the ego at time $t$ respectively. The following section explains how the spatiotemporal deep attention encoder processes this observation space. The encoder has two parts: a) Spatio-Temporal Attention Encoder (STAE) and b) Context Encoder (CE).
@@ IV-C Comparison Baselines
To thoroughly compare the proposed framework’s performance, the DAD-RL framework was evaluated against several baseline methods, and SOTA [11]. PPO: Proximal Policy Optimization stands out as a SOTA policy gradient approach with promising performance on robotic decision-making tasks. SAC: Soft Actor-Critic is another SOTA RL algorithm, an off-policy approach capitalizing on entropy maximization to assist agent training. DrQ: Data-regularized Q-learning employs a CNN-based augmentation to SAC, facilitating direct learning of a policy function from pixels. RDM: Rule-based Driver Model mimicking the driving behavior of neighboring vehicles within SMARTS simulations, offering a benchmark for evaluating the rule-based performance. DT: Decision Transformer is an innovative approach that reformulates the RL problem as a conditional sequence modeling task and employs a transformer to generate future actions from past states, actions, and rewards. SRT: Scene-Rep Transformer (SRT) is a transformer-based framework recently proposed in [11]. SRT employs two transformer architectures - a Multi-Stage Transformer (MST) for encoding the multi-modal scene input and another Sequential Latent Transformer (SRT) to instill predictive information into the latent state vector.